OpenStudy (anonymous):

A technical writer makes $45,000 per year and has just begun contributing 8% of her salary to her 401(k) plan. If her employer matches 6% of her contributions, how much will have been contributed to the 401(k) plan after 2 years?

OpenStudy (wolf1728):

salary = 45,000 8% of 45,000 = .08*45,000 = 3,600.00 employer 6%of 45,000 = 2,700.00 After one year the total amount contributed is 3,600 + 2,700 = 6,300.00 After 2 years, contribution = 2 * 6,300 = 12,600.00

OpenStudy (anonymous):

12,600.00 isn't one of the options